# Kramer’s Escape Rate and Phase Transition Dynamics in AdS Black Holes with Dark Structures

### v1

@@ 6 Conclusion
The study of black hole phase transitions based on the Gibbs free energy function is an important and influential part of advancing our understanding of black hole thermodynamics. However, a point that existed in this conventional method was that it seemed only static phase transitions were considered, leaving questions about the dynamics of phase transitions unanswered.
@@
The proposition that the dynamics of phase transitions in black holes are controlled by thermodynamic and statistical forces, followed by the study of phase transitions according to the stochastic Langevin equation or the Fokker-Planck equation [33], and considering the ’mean first passage time’ in some studies [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56], and then the introduction of Kramer’s escape rate in the study of phase transitions [60], motivated us to examine and compare different forms of the free energy landscape and then study Kramer’s escape rate in the first-order phase transitions of more complex black holes. These black holes are specifically formed based on cosmic and stringy properties, meaning they interact more with the features of dark structures. Our goal is to see what similarities or differences the addition of dark structure and stringy parameters create in the behavior of these black holes compared to the simpler model of the charged black hole.
According to this, in this article, we examined the escape rate from two perspectives. Initially, we considered it in terms of distance or the radius of the black hole and observed that black holes exhibit natural behavior consistent with environmental evidence. That is, the escape rate initially increases with the radius until it reaches a certain maximum, but this increase is not uncontrolled or permanent. A physical feedback mechanism causes this probability to gradually decrease, ending at a specific radius tolerable for the black hole’s survival. What is evident in comparison is that the escape rate and leakage from the small to the large black hole always dominate over the reverse. However,in the (NLM-C-Q-PFD) case, the escape rate diagram as a function of temperature (pressure) indicates something else. Meaning that the diagram initially, at its peak, and even near the end, shows behavior similar to the escape rate diagram as a function of distance. But suddenly, near the end, at a relatively high temperature (pressure), it takes on a different behavior. Specifically, at a certain temperature (pressure), the probability of transitioning from small to large and vice versa becomes equal. This implies that at this temperature, black holes have the ’choice’ to select their desired path towards either the small or large black hole. After this temperature, in a region, albeit small, the process completely reverses, meaning the probability of transitioning from the large to the small black hole becomes dominant, and black holes prefer to take the path from the large to the small black hole! But in the case of (EGB-YM-CS), after the intersection point of the graphs, the high probable transition still maintains its preference over the ’low probable’ transition and, in our opinion, shows a more perfect harmony with nature.We have addressed this in the discussion and comparison section. What we found in comparison with the charged black hole model is that in the charged model, this region is quite evident and noticeable. However, with the addition of parameters related to dark structure, this region-which we consider to be less physical-will either disappear (EGB-YM-CS) or become so small that radial changes due to their insignificance do not appear in the radius diagram, but still exist in the escape rate diagram (NLM-C-Q-PFD) as a function of temperature (pressure). The question of whether the existence of this region, albeit small, for a black hole has a physical interpretation and is considered natural behavior, and how it should be justified, is something that requires further study of various models. However, in conclusion, it seems that as the theoretical models of black holes evolve and more empirical and observable factors such as dark structure are added to the model, there will exist greater conformity with the ’pattern of repetition’ mentioned at the beginning of the introduction.

### v2

@@ 6 Conclusion
The study of black hole phase transitions based on the Gibbs free energy function has played a significant and influential role in advancing our understanding of black hole thermodynamics. However, a common issue with this conventional approach is that it seems to consider phase transitions only in a static form, leaving questions about the dynamics of phase transitions largely unanswered. The proposition that the dynamics of phase transitions in black holes are governed by thermodynamic and statistical forces has led to the possibility of examining phase transitions based on the stochastic Langevin equation or the Fokker-Planck equation [33], considering the ’mean first passage time’[34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. In this context, the introduction of the Kramers escape rate in the study of first-order phase transitions[60] prompted us to examine various free energy landscape forms and subsequently investigate the Kramers escape rate for the phase transitions of more complex black holes . These black holes are specifically formed based on cosmic and stringy features, meaning they interact more with the characteristics of dark structures. Our goal was to see what effect the addition of dark structure and stringy parameters would have on the dynamic behavior of phase transitions in this category of black holes.
Accordingly, in this paper, we have dynamically investigated the transition from a small black hole to a large black hole in first-order phase transitions from the perspective of the Kramers escape rate. Considering the escape rate as a function of distance or black hole radius, we examined whether the dynamic behavior of this phase transition for the black hole under study is exactly as expected at all moments of the transition. That is, firstly, the escape rate initially increases with the radius from a local minimum towards a global minimum and, after reaching a certain maximum, gradually decreases due to the activation of a reaction mechanism until it tends to zero. Secondly, the rate of escape and leakage from a small black hole to a large black hole always follows a logical and direct trend.
With this background, we studied the models (NLM-C-Q-PFD) and (EGB-YM-CS) and found that the escape rate diagram as a function of radius shows something different in specific regions. This means that the diagrams at the beginning, middle, and near the end of phase transition, at a specific radius, show equal probabilities of transition from a small to a large black hole and vice versa, which we have called contact coordinates in this paper. Beyond this radius, in a specific range up to the final point of phase transition, the reverse process becomes completely probable, meaning the probability of transitioning from a large black hole to a small black hole becomes dominant, and it seems that black holes prefer to travel from a large black hole to a small one. An interesting point about this region is that at the onset of phase transition, this area is very small, and as we approach the end in the sequence of frames, this area grows significantly.
Since, on the one hand, the growth of the probability of the reverse process corresponds to the increase in radius and, on the other hand, it is in the direction of the end of the phase transition process, it seems that in fact, this reverse trend is a reactive mechanism by the black hole. It is actually an effort to limit and control the phase transition to prevent excessive radial growth and prevent any uncontrolled instability so that the black hole can conclude the process towards a larger black hole under conditions that preserve the stability of the final structure and do not disrupt the survival of the black hole.